# ICP-4D: Bridging Iterative Closest Point and LiDAR Panoptic Segmentation

### v1

@@ Abstract
Dominant paradigms for 4D LiDAR panoptic segmentation are usually required to train deep neural networks with large superimposed point clouds or design dedicated modules for instance association. However, these approaches perform redundant point processing and consequently become computationally expensive, yet still overlook the rich geometric priors inherently provided by raw point clouds. To this end, we introduce ICP-4D, a simple yet effective training-free framework that unifies spatial and temporal reasoning through geometric relations among instance-level point sets. Specifically, we apply the Iterative Closest Point (ICP) algorithm to directly associate temporally consistent instances by aligning the source and target point sets through the estimated transformation. To stabilize association under noisy instance predictions, we introduce a Sinkhorn-based soft matching. This exploits the underlying instance distribution to obtain accurate point-wise correspondences, resulting in robust geometric alignment. Furthermore, our carefully designed pipeline, which considers three instance types—static, dynamic, and missing—offers computational efficiency and occlusion-aware matching. Our extensive experiments across both SemanticKITTI and panoptic nuScenes demonstrate that our method consistently outperforms state-of-the-art approaches, even without additional training or extra point cloud inputs.
**footnotetext: Corresponding author.
@@ 1 Introduction
Autonomous driving systems heavily depend on precise perception of dynamic 3D environments with geometrically accurate measurements captured by LiDAR sensors. In particular, LiDAR segmentation plays a crucial role in enabling comprehensive point-wise understanding. In 3D environments, LiDAR panoptic segmentation [19, 32] originates from semantic and instance segmentation, which jointly identify class categories and instance IDs. Along with the recent significant advances in frame-wise perception, 4D LiDAR panoptic segmentation [2] has emerged to capture temporal relationships in a holistic manner. In this context, ensuring temporal continuity is a principal purpose in 4D LiDAR panoptic segmentation, even under instance motions.
@@
With its potential benefits, as shown in Figure 1, the mainstream of prior works can be categorized into three discrete paradigms: 1) IoU-based association, 2) Query-propagated association, and 3) Detect & Track association. One solution [2, 20, 63, 54, 16], superimposing point clouds, is widely adopted for establishing 4D spatio-temporal volume. It enables consistent instance tracking over time by measuring the overlapped regions (see Figure 1(a)). However, despite its simplicity, stacking point clouds (i.e., increasing the number of points per batch) leads to a quadratic growth in computational overhead, which is an inherent challenge in point cloud processing [28, 22, 50, 51].
@@
An alternative direction [30, 1, 29] seeks to mitigate this issue by treating each scan independently, as illustrated in Figure 1(b) and 1(c). They are primarily derived from the two major paradigms of multi-object tracking, tracking-by-detection [7, 49, 58, 36] and query propagation [55, 59, 31], to assign consistent IDs for instances. Both approaches differ substantially in training strategies, yet commonly suffer from error accumulation [59, 41] and reliance on costly and dataset-dependent training. Consequently, we argue that this suboptimal design fails to fully exploit the strong perception capability of 3D panoptic networks, which naturally raises the following question: Can instance association be achieved efficiently using only a 3D panoptic network?
@@
Figure 1: Comparison of 4D LiDAR panoptic segmentation methods: (a) IoU-based, (b) Query-propagated, (c) Detect & Track, and (d) ICP-4D. Instead of relying on training with large-scale point clouds, ours achieves reliable association in a fully training-free manner. Refer to caption indicates methods that require training and Refer to caption denotes frozen network.
To answer this, we propose Iterative Closest Point (ICP)-based strategy for reliable instance matching, even without any training process. ICP is widely used for point cloud registration [46, 57, 11], but its applicability beyond that has been largely underexplored. Here, we first integrate ICP into the 4D LiDAR panoptic segmentation framework. This allows explicit mapping through estimated transformation between instance segment pairs in successive scans. While this mapping works well with complete segment pairs, noisy predictions hinder finding corresponding points via nearest-neighbor approach, as it ignores the global instance structure. To handle this, we propose Sinkhorn-based soft matching that strengthens the robustness against outliers with instance-aware correspondences. Specifically, we represent each instance’s point set as a probability distribution. Then, the correspondence problem is cast as a Sinkhorn Optimal Transport (OT) problem [12], which computes an entropy-regularized transport plan between the two distributions. This plan identifies point pairs that minimize the transportation cost, thereby establishing soft correspondences.
ICP demonstrates competitive performance in instance association. However, its iterative mechanism and the physical limitations of LiDAR sensors introduce remaining challenges, a marginal increase in runtime and vulnerability to occlusion. Given these limitations, we classify instances into three categories: static, dynamic, and missing. For associating static instances, we employ its point set statistics (i.e., mean and covariance) considering the rigid property. This significantly reduces the number of matching candidates and preserves computational efficiency. Moreover, to compensate for occluded instances, we present a memory bank to store prediction results within a short temporal window.
@@
To wrap it up, we named our overall pipeline as ICP-4D and demonstrate its effectiveness through extensive experiments on the SemanticKITTI [4] and nuScenes [13] benchmarks. Note that ICP-4D achieves superior performance over existing state-of-the-art methods and ranks 1st on the SemanticKITTI test leaderboard, even under the minimal setting of a single scan and without any additional training.
@@
  • We introduce ICP-4D that is a novel training-free 4D LiDAR panoptic segmentation pipeline, flexibly integrated with a wide range of 3D panoptic segmentation networks.
  • We propose a Sinkhorn-based soft matching strategy to enhance the robustness of point correspondences, and we delicately design instance state-conditioned solutions to effectively mitigate time complexity and occlusion.
  • Extensive experiments in challenging environments clearly demonstrate the superiority of ICP-4D, supported by in-depth analysis.
@@ 2.1 LiDAR Panoptic Segmentation
LiDAR panoptic segmentation simultaneously handles two different basis tasks, semantic and instance segmentation. In early, proposal-based methods [27, 34, 52] filter the background points with Mask R-CNN [14]’s semantic head, then group the thing points into instances. In contrast, proposal-free methods [15, 37, 62, 25] independently predict semantic prediction and class-agnostic instance segments via end-to-end training manners. 4D LiDAR panoptic segmentation extends these frameworks to sequential scans, requiring instance IDs to remain consistent over time.
@@
In terms of association strategy, predominant works [60, 2, 20, 63, 54, 16] aggregate consecutive point clouds (e.g., 2-4 scans) and group instances in overlapping regions to ensure temporal continuity. Alternatively, CA-Net [29] associates instances via feature-wise similarity, followed by the outputs of 3D panoptic network. Mask4D [30] preserves identity consistency across scans by iteratively updating the track queries in an end-to-end manner. Despite these attempts, we observe significant issues regarding the computational cost and limited adaptability (see Sec. 1). Instead, we propose ICP-4D that is a flexible, efficient, and well-balanced instance association through direct geometric registration.
2.2 Multi-Object Tracking (MOT)
@@
MOT is an indispensable task for autonomous driving systems, enabling the tracking of all detected objects across scans. Tracking-by-detection (TBD) paradigm [7, 49, 58, 36] has become the de facto standard due to its simplicity and the rapid progress of detectors. Here, the motion model (e.g., Kalman Filter [18]) plays a key role by estimating future object states to maintain tracking continuity. The emergence of DETR [64] has recently drawn increasing attention to the end-to-end paradigm [55, 59, 31], which performs detection and tracking in parallel using query-based tracking. These paradigms are not confined to 2D images but are increasingly embraced in 3D autonomous driving scenarios [48, 61, 35, 56] as well. However, despite sharing objectives with MOT, 4D LiDAR panoptic segmentation differs fundamentally, as it requires holistic point-level scene understanding instead of bounding boxes. Although ICP-4D follows the TBD paradigm, it operates directly on point-level panoptic outputs, enabling instance association grounded in the geometric structure of the underlying point clouds.
2.3 Iterative Closest Point (ICP)
@@
ICP [6] is one of the most well-established and widely adopted algorithm for point-set registration [9, 24, 39]. Over the years, many variants of the vanilla ICP have been proposed to address its limitations, and have been widely adopted in both traditional engineering [8, 38, 40, 53] and deep learning–based applications [46, 57, 11]. Nowadays, ICP-Flow [26] proposes a compelling scene flow prediction framework that inspired this work. They only consider point-wise motion estimation between two temporally adjacent scans by leveraging well-separated instance clusters. In this paper, we introduce ICP into the context of the 4D LiDAR panoptic segmentation, ensuring the temporal continuity of instances across the entire scans. Remarkably, our Sinkhorn-based soft matching enables instance-centric association, even when predictions are noisy, imprecise, or fragmented.
@@ 3 Method
ICP-4D primarily focuses on ensuring temporal continuity of the instances in an unsupervised manner. To handle this, we adopt Iterative Closest Point (ICP) [6] to associate instances through their geometric consistency over time. In this section, we first provide the introductory ICP (Sec. 3.1), then go over the technical details of our methods (Sec. 3.2).
@@ 3.1 Preliminary
Iterative Closest Point (ICP). Conventionally, for point cloud registration, ICP estimates the optimal transformation by iteratively minimizing the difference between source and target point sets. Given two point sets $\mathcal{P}_{src}\in\mathbb{R}^{n\times 3}$ and $\mathcal{P}_{dst}\in\mathbb{R}^{m\times 3}$, it starts with the initial guess of rigid transformation ($\mathbf{R}^{(0)},\mathrm{t}^{(0)}$), where the rotation $\mathbf{R}\in\text{SO}(3)$ and translation $\mathrm{t}\in\mathbb{R}^{3}$, which satisfies the following formula:
$\mathcal{P}_{dst}\approx\mathbf{R}\mathcal{P}_{src}+\mathrm{t}\;.$ (1)
Here, in order to compute the alignment error, pairwise correspondences are established using a nearest-neighbor search algorithm (e.g., KD-tree [5]). Classical ICP adopts $l_{2}$ distance as the error metric between the matched point pairs. Ultimately, the optimal rigid transformation is obtained by iteratively solving a series of least-squares problems as:
@@
to update the rotation $\mathbf{R}$ and translation $\mathrm{t}$ until convergence. With this theoretical background, we investigate how ICP contributes to consistent instance ID assignment across frames while enhancing robustness against outliers.
3.2 ICP-4D
@@
Our proposed ICP-4D is developed on top of a pre-trained 3D LiDAR panoptic segmentation model [15, 54], which is specialized in frame-wise perception. Through this model, we obtain the per-point instance and semantic prediction $(\mathcal{I}_{i},\mathcal{S}_{i})$ for a given point cloud with $N$ points, where $\mathcal{S}_{i}\in\{1,...,C\}$, $\mathcal{I}_{i}\in\mathbb{N}_{0}$, and $C$ is the number of classes. Prior to the association stage, we construct a set of matching candidates between $t$ and t1t-1 times by pairing source and destination instances as follows:
$\mathcal{G}_{\text{src}}\times\mathcal{G}_{\text{dst}}=\{(a_{i},b_{j})\;|\;a_{i}\in\mathcal{G}_{\text{src}},\;b_{j}\in\mathcal{G}_{\text{dst}}\},$ (3)
@@
where $\mathcal{G}_{\text{src}}$ and $\mathcal{G}_{\text{dst}}$ are respectively denoted as the sets of unique instance IDs in the source and destination. Then, we derive the filtered set $\mathcal{M}=\mathcal{G}^{\prime}_{\text{src}}\times\mathcal{G}^{\prime}_{\text{dst}}$ by keeping only the pairs in $\mathcal{G}_{\text{src}}\times\mathcal{G}_{\text{dst}}$ that share the same semantic class. In this study, we categorize instances into dynamic, static, and missing types to facilitate effective instance association across frames.
3.2.1 Static Instance
@@
Under the assumption that both the center and the shape of a non-moving rigid object remain invariant, we filter static instances based on their centers and further refine them using covariance cues to improve selection fidelity. Given the matching candidates $\mathcal{M}$, we first represent each instance by the set of 3D points belonging to it. For all instances in $\mathcal{G}^{\prime}_{\text{src}}$ and $\mathcal{G}^{\prime}_{\text{dst}}$, we denote their corresponding point sets as $P(\mathcal{G}^{\prime}_{\text{src}})$ and $P(\mathcal{G}^{\prime}_{\text{dst}})$, respectively. Then, we compute the center of each point set:
$\mu_{s}={1\over|P(\mathcal{G}^{\prime}_{\text{src}})|}\sum_{u\in P(\mathcal{G}^{\prime}_{\text{src}})}\!\!\!u,\quad\mu_{d}={1\over|P(\mathcal{G}^{\prime}_{\text{dst}})|}\sum_{v\in P(\mathcal{G}^{\prime}_{\text{dst}})}\!\!\!v,$ (4)
@@
where $u,v\in\mathbb{R}^{3}$ and indicate the world coordinates of points. We keep only the pairs satisfying $\mu_{d}-\mu_{s}||_{2}<\tau_{center}$, where $\tau_{center}$ is the distance threshold for identifying non-moving objects based on center consistency. Although this naive approach performs well under ideal conditions, it suffers when applied to noisy predicted point sets, resulting in unsuitable pair associations. Thus, we introduce an additional covariance-based constraint that measures the similarity of the spatial spread and geometric structure of the instance point sets. Specifically, the covariance of each point set is computed as follows:
${\sigma}_{s}=\mathrm{Cov}(P(\mathcal{G}^{\prime}_{\text{src}})),\quad{\sigma}_{d}=\mathrm{Cov}(P(\mathcal{G}^{\prime}_{\text{dst}})).$ (5)
@@
where $\text{tr}(\cdot)$ and $\|\cdot\|_{F}$ denote the matrix trace and the Frobenius norm. Then, based on the obtained similarity scores, we choose the highest score for each source instance to find the best-matching pairs while avoiding duplicate associations.
Figure 2: Illustration of Sinkhorn-based soft matching. (a) Nearest neighbor-based matching aligns each source point with its closest counterpart, focusing only on local point proximity. (b) In contrast, correspondences are computed by transport plan $\mathcal{Q}$, enabling instance-aware matching that respects the global geometry of each instance point set.
3.2.2 Dynamic Instance
After filtering static instances, the remaining candidates $\mathcal{G}_{\text{src}}^{\prime\prime}$ and $\mathcal{G}_{\text{dst}}^{\prime\prime}$ are fed into the ICP loop that is the core mechanism of our framework that enables instance association without any supervised training. In the vanilla ICP loop, the first step is finding point pairs to compute the distance by utilizing the nearest-neighbor algorithm [5]. While ICP has a powerful ability of registering point cloud sets, determining the corresponding points under noisy instance predictions potentially causing erroneous matches. Therefore, we propose a Sinkhorn-based soft matching scheme that facilitates stable correspondence estimation within the ICP framework.
Sinkhorn-based Soft Matching. Sinkhorn-based soft matching starts from regarding the point cloud as a probability distribution rather than discrete set of individual points. Given two point sets $P(\mathcal{G}_{\text{src}}^{\prime\prime})=\{u_{i}\}^{I-1}_{0}$ and $P(\mathcal{G}_{\text{dst}}^{\prime\prime})=\{v_{j}\}^{J-1}_{0}$, where $I$ and $J$ denote the number of points in each segment, we formulate correspondence estimation as an Optimal Transport (OT) problem [12]. Specifically, we seek a transport plan $\mathcal{Q}$ that minimizes the cost of transforming one distribution into the other as follows:
$\text{OT}(P(\mathcal{G}^{\prime\prime}_{\text{src}}),\;P(\mathcal{G}^{\prime\prime}_{\text{dst}})):=\underset{\mathcal{Q}}{\operatorname*{min}}\langle\mathcal{Z},\;\mathcal{Q}\rangle_{F}-\epsilon H(\mathcal{Q}),$ (7)

### v2

# Training-Free Global Geometric Association for 4D LiDAR Panoptic Segmentation

@@ Abstract
Dominant paradigms for 4D LiDAR panoptic segmentation are usually required to train deep neural networks with large superimposed point clouds or design dedicated modules for instance association. However, these approaches perform redundant point processing and consequently become computationally expensive, yet still overlook the rich geometric priors inherently provided by raw point clouds. To this end, we introduce Geo-4D, a simple yet effective training-free framework that unifies spatial and temporal reasoning, enabling holistic LiDAR perception over long time horizons. Specifically, we propose a global geometric association strategy that establishes consistent instance correspondences by estimating an optimal transformation between instance-level point sets. To mitigate instability caused by structural inconsistencies in point cloud observations, we propose a global geometry-aware soft matching mechanism that enforces spatially coherent point-wise correspondences grounded in the spatial distribution of instance point sets. Furthermore, our carefully designed pipeline, which considers three instance types—static, dynamic, and missing—offers computational efficiency and occlusion-aware matching. Our extensive experiments across both SemanticKITTI and nuScenes demonstrate that our method consistently outperforms state-of-the-art approaches, even without additional training or extra point cloud inputs.
@@ 1 Introduction
Autonomous driving systems heavily depend on precise perception of dynamic 3D environments with geometrically accurate measurements captured by LiDAR sensors. In particular, LiDAR segmentation plays a crucial role in enabling comprehensive point-wise understanding. In 3D environments, LiDAR panoptic segmentation [kirillov2019panoptic, milioto2020lidar] originates from semantic and instance segmentation, which jointly identify class categories and instance IDs. Along with the recent significant advances in frame-wise perception, 4D LiDAR panoptic segmentation [Aygun_2021_CVPR] has emerged to capture temporal relationships in a holistic manner. In this context, ensuring temporal continuity is a principal purpose in 4D LiDAR panoptic segmentation, even under instance motions.
@@
With its potential benefits, as shown in Figure 1, the mainstream of prior works can be categorized into three discrete paradigms: 1) IoU-based association, 2) Query-propagated association, and 3) Detect & Track association. One solution [Aygun_2021_CVPR, kreuzberg20224d, zhu20234d, yilmaz2024mask4former, 10380457], superimposing point clouds, is widely adopted for establishing 4D spatio-temporal volume. It enables consistent instance tracking over time by measuring overlapping regions (see Figure 1(a)). However, despite its simplicity, stacking point clouds (i.e., increasing the number of points per batch) leads to a quadratic growth in computational overhead, which is an inherent challenge in point cloud processing [liu2023flatformer, lai2022stratified, wu2022point, wu2024point].
@@
An alternative direction [marcuzzi2023mask4d, athar20234d, marcuzzi2022contrastive] seeks to mitigate this issue by treating each scan independently, as illustrated in Figure 1(b) and 1(c). They are primarily derived from the two major paradigms of multi-object tracking, tracking-by-detection [bewley2016simple, wojke2017simple, zhang2022bytetrack, qin2024towards] and query propagation [zeng2022motr, zhang2023motrv2, meinhardt2022trackformer], to assign consistent IDs for instances. Both approaches differ substantially in training strategies, yet commonly suffer from error accumulation [zhang2023motrv2, segu2024samba] and reliance on costly and dataset-dependent training. Consequently, we argue that this suboptimal design fails to fully exploit the strong perceptual capabilities of 3D panoptic networks, which naturally raises the following question: Can instance association be achieved efficiently using only a 3D panoptic network?
@@
Figure 1: Comparison of 4D LiDAR panoptic segmentation methods: (a) IoU-based, (b) Query-propagated, (c) Detect & Track, and (d) Geo-4D. Instead of relying on training with large-scale point clouds, ours achieves reliable association in a fully training-free manner. Refer to caption indicates methods that require training and Refer to caption denotes frozen network.
To answer this, we propose a global geometric association strategy for consistent and reliable instance correspondence, even without any training process. Iterative Closest Point (ICP) is widely used for point cloud registration [wang2019deep, zhang20233d, choy2020deep], but its applicability beyond that has been largely underexplored. Here, we integrate ICP into the 4D LiDAR panoptic segmentation framework, thereby enabling explicit instance correspondence through geometric alignment under estimated inter-scan transformations. Additionally, we introduce a novel paradigm for point-wise correspondence, Global Geometry-aware Soft Matching (GGSM), enhancing the robustness to structural noise by enforcing instance-aware geometric correspondences. Traditional ICP performs well on point cloud pairs under ideal conditions, but it relies on nearest-neighbor point correspondences. In real-world noisy measurements acquired from on-vehicle sensors, such point-wise matching often yields incorrect transformations and ultimately leads to erroneous instance associations, as it ignores the global instance structure. In contrast, our GGSM allows reliable correspondences based on the overall geometric context rather than local proximity. Specifically, it identifies point pairs that minimize the transport cost by solving the earth mover’s problem [cuturi2013sinkhorn] between two instance point set distributions. As a result, instance association shifts from greedy local pairing to globally consistent alignment, yielding more stable and accurate matching even under challenging environments.
In addition, in real-world observations, instances may remain stationary, move across frames, or temporarily disappear due to occlusion or sensing limitations. Motivated by these instance states, we classify instances into three types: static, dynamic, and missing. For static instances, we apply statistics-based geometric filtering using the mean and covariance of their point distributions to preserve spatial consistency across frames. This filtering step also significantly reduces the number of matching candidates before applying global geometric association to dynamic instances, thereby ensuring computational efficiency. Moreover, we incorporate a short-term memory bank to recover temporarily unobservable instances. We retain only unmatched instances within a short-term window, preserving temporal consistency with minimal memory overhead.
@@
To wrap it up, we named our overall pipeline as Geo-4D and demonstrate its effectiveness through extensive experiments on the SemanticKITTI [behley2019semantickitti] and nuScenes [fong2022panoptic] benchmarks. Note that Geo-4D achieves superior performance over existing state-of-the-art methods and ranks 1st on the SemanticKITTI test leaderboard, even under the minimal setting of a single scan and without any additional training.
@@
  • We introduce Geo-4D that is a novel training-free 4D LiDAR panoptic segmentation pipeline, flexibly integrated with a wide range of 3D panoptic segmentation networks.
  • We propose a global geometric association strategy to enhance the robustness of instance correspondences, and we delicately design instance state-conditioned solutions to effectively mitigate time complexity and occlusion.
  • Extensive experiments in challenging environments clearly demonstrate the superiority of Geo-4D, supported by in-depth analysis.
@@ 2 Related Work
LiDAR Panoptic Segmentation. LiDAR panoptic segmentation simultaneously handles two different basis tasks, semantic and instance segmentation. In early, proposal-based methods [liu2019end, mohan2021efficientps, xiong2019upsnet] filter the background points with Mask R-CNN [he2017mask]’s semantic head, then group the thing points into instances. In contrast, proposal-free methods [hong2021lidar, razani2021gp, zhou2021panopticp, li2022panoptic] independently predict semantic prediction and class-agnostic instance segments via end-to-end training manners. 4D LiDAR panoptic segmentation extends these frameworks to sequential scans, requiring instance IDs to remain consistent over time.
@@
In terms of association strategy, predominant works [zhang2025zero, Aygun_2021_CVPR, kreuzberg20224d, zhu20234d, yilmaz2024mask4former, 10380457] aggregate consecutive point clouds (e.g., 2-4 scans) and group instances in overlapping regions to ensure temporal continuity. Alternatively, CA-Net [marcuzzi2022contrastive] associates instances via feature-wise similarity, followed by the outputs of 3D panoptic network. Mask4D [marcuzzi2023mask4d] preserves identity consistency across scans by iteratively updating the track queries in an end-to-end manner. Despite these attempts, we observe significant issues regarding the computational cost and limited adaptability (see Sec. 1). Instead, we propose Geo-4D that is a flexible, efficient, and well-balanced instance association through direct geometric registration.
@@
Multi-Object Tracking (MOT). MOT is an indispensable task for autonomous driving systems, enabling the tracking of all detected objects across scans. Tracking-by-detection (TBD) paradigm [bewley2016simple, wojke2017simple, zhang2022bytetrack, qin2024towards] has become the de facto standard due to its simplicity and the rapid progress of detectors. Here, the motion model (e.g., Kalman Filter [kalman1960new]) plays a key role by estimating future object states to maintain tracking continuity. The emergence of DETR [zhu2020deformable] has recently drawn increasing attention to the end-to-end paradigm [zeng2022motr, zhang2023motrv2, meinhardt2022trackformer], which performs detection and tracking in parallel using query-based tracking. These paradigms are not confined to 2D images but are increasingly embraced in 3D autonomous driving scenarios [weng2020ab3dmot, zhou2020tracking, pang2022simpletrack, zhang2022mutr3d] as well. However, despite sharing objectives with MOT, 4D LiDAR panoptic segmentation differs fundamentally, as it requires holistic point-level scene understanding instead of bounding boxes. Although Geo-4D follows the TBD paradigm, it operates directly on point-level panoptic outputs, enabling instance association grounded in the geometric structure of the underlying point clouds.
@@
Iterative Closest Point (ICP). ICP [besl1992method] is one of the most well-established and widely adopted algorithms for point-set registration [bustos2017guaranteed, li2021practical, rusu2009fast]. Over the years, many variants of the vanilla ICP have been proposed to address its limitations, and have been widely adopted in both traditional engineering [bouaziz2013sparse, rusinkiewicz2001efficient, segal2009generalized, yang2015go] and deep learning–based applications [wang2019deep, zhang20233d, choy2020deep]. Nowadays, ICP-Flow [lin2024icp] proposes a compelling scene flow prediction framework that inspired this work. They only consider point-wise motion estimation between two temporally adjacent scans by leveraging well-separated instance clusters. In this paper, we introduce ICP into the context of the 4D LiDAR panoptic segmentation, ensuring the temporal continuity of instances across the entire scans. Remarkably, our Global Geometry-aware Soft Matching enables instance-centric association, even when instance point sets are noisy, imprecise, or fragmented.
@@ 3 Method
Geo-4D primarily focuses on ensuring temporal continuity of the instances in a training-free manner. To handle this, we adopt Iterative Closest Point (ICP) [besl1992method] to associate instances through their geometric consistency over time. In this section, we first provide a brief introduction to ICP (Sec. 3.1), then go over the technical details of our methods (Sec. 3.2).
@@ 3.1 Preliminary
Iterative Closest Point (ICP). Conventionally, for point cloud registration, ICP estimates the optimal transformation by iteratively minimizing the difference between source and target point sets. Given two point sets $\mathcal{P}_{\text{src}}=\{p_{1},...,p_{I}\}$ and $\mathcal{P}_{\text{dst}}=\{q_{1},...,q_{J}\}$ in $\mathbb{R}^{3}$, it starts with the initial guess of rigid transformation ($\mathbf{R}^{(0)},\mathbf{t}^{(0)}$), where the rotation $\mathbf{R}\in\text{SO}(3)$ and translation $\mathbf{t}\in\mathbb{R}^{3}$, which satisfies the following formula:
$q\approx\mathbf{R}p+\mathbf{t}\;.$ (1)
Here, in order to compute the alignment error, pairwise correspondences are established using a nearest-neighbor search algorithm (e.g., KD-tree [bentley1975multidimensional]) as:
@@
where $k$ denotes the iteration. Classical ICP adopts $l_{2}$ distance as the error metric between the matched point pairs. Ultimately, the optimal rigid transformation is obtained by iteratively solving a series of least-squares problems as:
[…]
@@
to update the rotation $\mathbf{R}$ and translation $\mathbf{t}$ until convergence. With this theoretical background, we investigate how ICP contributes to consistent instance ID assignment across frames while enhancing robustness against outliers.
3.2 Geo-4D
@@
Our proposed Geo-4D is developed on top of a pre-trained 3D LiDAR panoptic segmentation model [hong2021lidar, yilmaz2024mask4former], which is specialized in frame-wise perception. Following the standard assumption in LiDAR panoptic segmentation, we transform all points into the world coordinate system using the ego pose. Through this model, we obtain the per-point semantic and instance prediction $(s_{i},\;\ell_{i})$ for a given point cloud, where $s_{i}\in\{1,...,S\}$, $\ell_{i}\in\mathbb{N}_{0}$, and $S$ is the number of classes. Prior to the association stage, we construct a set of matching candidates between $t$ and t1t-1 times by pairing source and destination instances as follows:
$\mathcal{M}=\mathcal{G}_{\text{src}}\times\mathcal{G}_{\text{dst}}=\{(a,b)\;|\;a\in\mathcal{G}_{\text{src}},\;b\in\mathcal{G}_{\text{dst}}\},$ (4)
@@
where $\mathcal{G}_{\text{src}}$ and $\mathcal{G}_{\text{dst}}$ are respectively denoted as the sets of unique instance IDs in the source and destination. Then, we derive the filtered set $\mathcal{M}^{\prime}$ by keeping only the pairs in $\mathcal{M}$ that share the same semantic class. In this study, we categorize instances into dynamic, static, and missing types to facilitate effective instance association across frames.
@@
Static Instance. Under the assumption that both the center and the shape of a non-moving rigid object remain invariant, we filter static instances based on their centers and further refine them using covariance cues to improve selection fidelity. Given the matching candidates $\mathcal{M}^{\prime}$, we first represent each instance by the set of 3D points belonging to it. For each instance pair $(a_{\text{src}}^{\prime},\;b_{\text{dst}}^{\prime})\in\mathcal{M}^{\prime}$, we denote the corresponding point sets as $\mathcal{P}(a_{\text{src}}^{\prime})$ and $\mathcal{P}(b_{\text{dst}}^{\prime})$, respectively. Then, we compute the centroid of each point set:
$\mu_{s}=\frac{1}{|\mathcal{P}(a^{\prime}_{\text{src}})|}\sum_{p\in\mathcal{P}(a^{\prime}_{\text{src}})}p,\qquad\mu_{d}=\frac{1}{|\mathcal{P}(b^{\prime}_{\text{dst}})|}\sum_{q\in\mathcal{P}(b^{\prime}_{\text{dst}})}q,$ (5)
@@
where $p,q\in\mathbb{R}^{3}$ represent the world coordinates of points. We keep only the pairs satisfying $\mu_{d}-\mu_{s}||_{2}<\tau_{center}$, where $\tau_{center}$ is the distance threshold for identifying non-moving objects based on center consistency. Although this naive approach performs well under ideal conditions, it suffers when applied to noisy predicted point sets, resulting in unsuitable pair associations. Thus, we introduce an additional covariance-based constraint that measures the similarity of the spatial spread and geometric structure of the instance point sets. Specifically, the covariance of each point set is computed as follows:
${\sigma}_{s}=\mathrm{Cov}(\mathcal{P}(a^{\prime}_{\text{src}})),\quad{\sigma}_{d}=\mathrm{Cov}(\mathcal{P}(b^{\prime}_{\text{dst}})).$ (6)
@@
where $\text{tr}(\cdot)$ and $\|\cdot\|_{F}$ denote the matrix trace and the Frobenius norm. Then, for all candidate pairs satisfying the filtering conditions, we choose the smallest discrepancy in Eq. 7 for each source instance to determine the best-matching pairs while avoiding duplicate associations.
Figure 2: (a) Nearest neighbor aligns each source point with its closest counterpart, focusing only on local point proximity. (b) In contrast, our Global-Geometry-aware Soft Matching computes correspondences via an optimal transport plan $\hat{\mathcal{Q}}$, enabling instance-aware matching that respects the global geometry of each instance point set.
Dynamic Instance. After filtering static instances, we construct the remaining candidate pair set $\mathcal{M}^{\prime\prime}$. These candidate pairs are then fed into the global geometric association, which serves as the core component of our framework for training-free instance association. In the vanilla ICP loop, the first step is finding point pairs to compute the distance by utilizing the nearest-neighbor algorithm [bentley1975multidimensional]. While ICP has a powerful ability to register point cloud sets, determining the corresponding points under noisy instance predictions potentially causes erroneous matches. Therefore, we propose a Global Geometry-aware Soft Matching that facilitates stable correspondence estimation within the ICP.
Global Geometry-aware Soft Matching starts by regarding the point cloud as a probability distribution rather than a discrete set of individual points. For a candidate pair $(a^{\prime\prime}_{\text{src}},\;b^{\prime\prime}_{\text{dst}})\in\mathcal{M}^{\prime\prime}$, we denote their corresponding point sets as $\mathcal{P}(a^{\prime\prime}_{\text{src}})=\{p_{i}\}_{0}^{I-1}$ and $\mathcal{P}(b^{\prime\prime}_{\text{dst}})=\{q_{j}\}_{0}^{J-1}$, where $I$ and $J$ denote the number of points in each instance. We then formulate the correspondence estimation as an Optimal Transport (OT) problem [cuturi2013sinkhorn]. Specifically, we seek a transport plan $\mathcal{Q}$ that minimizes the cost of transforming one distribution into the other as follows:
$\text{OT}(\mathcal{P}(a^{\prime\prime}_{\text{src}}),\;\mathcal{P}(b^{\prime\prime}_{\text{dst}})):=\underset{\mathcal{Q}}{\operatorname*{min}}\langle\mathcal{Z},\;\mathcal{Q}\rangle_{F}-\epsilon H(\mathcal{Q}),$ (8)